Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of theOne-third supervisory salaries, two-thirds depreciation of special equipment (no resale value) Required: 1. Assuming the com2. Should the outside suppliers offer be accepted? 3. Suppose that if the carburetors were purchased, Troy Engines, Ltd., coRequired: 1. Assuming the company has no alternative use for the facilities that are now being used to produce the carburetorTuvuruye uiguuvarray Urvuyilly 17,000 LdluuleLUIS HU he vulsiue suppler? 2. Should the outside suppliers offer be accepted?
Answer #1

1) Differential analysis

Make Buy
Direct material 238000
Direct labor 136000
variable manufacturing overhead 51000
Fixed manufacturing overhead (51000/3) 17000
Purchase cost (17000*32) 544000
Total relevant cost 442000 544000

Financial (disadvantage) = 442000-544000 = -102000

2) No

3) Differential analysis

Make Buy
Direct material 238000
Direct labor 136000
variable manufacturing overhead 51000
Fixed manufacturing overhead (51000/3) 17000
Opportunity Cost 170000
Purchase cost (17000*32) 544000
Total relevant cost 612000 544000

Financial advantage = 612000-544000 = 68000

4) Yes